# SPECIAL RELATIVITY

### v11

@@ 16.8 Mathematical Analysis of Relativistic Velocity Composition
As a consequence of the space-time geometry, Lorentz coordinate systems are connected by Lorentz transformations. Consider a four-vector, $x=(x_{0},\vec{x})$, with squared-length $x^{2}=\eta_{ij}x^{i}x^{j}=(x_{0})^{2}-|\vec{x}|^{2}$ where $\eta_{ij}=\mathrm{diag}[1,-1,-1,-1]$ is the Minkowski space-time metric. Since Lorentz transformations preserve the space-time interval, they satisfy $x^{\prime}=\Lambda x$ such that $\eta_{ij}=\Lambda^{m}_{i}\Lambda^{n}_{j}\eta_{mn}$. This equation defines the most general 4 × 4 Lorentz transformation matrix $\Lambda$. The set of all such matrices forms a symmetry-group of the Minkowski metric under matrix multiplication, denoted by $O(1,3)$. In particular, $O(1,3)$ is a Lie group, appropriately called the Lorentz group.
@@
It is widely accepted that relationship between inertial frames are completely determined by their relative velocities and relative orientations. Consequently, the generic Lorentz transformation denoted as $\Lambda(\vec{v},\vec{\theta})$, is parameterized by the velocity $\vec{v}$ and the rotation parameter $\vec{\theta}$, representing a simultaneous boost and rotation. Here, $\vec{v}$ is a vector in three-dimensional Euclidean space, wile $\vec{\theta}$ belongs to the group $O(3)$, which consists of all possible spatial rotations in three-dimensions. This parameterization of the Lorentz group is commonly found in textbooks.
@@
Now, we turn to a significant challenge associated with this parametrization of the Lorentz group. A comparison with the Galilean group may provide insight. It is well known that the Galilean transformation $G(\vec{v},\vec{\theta})$, which preserves the group composition law, is isomorphic to the semi-direct product group $R^{3}\times O(3)$, where $R^{3}$ represents the group of Galilean boosts and $O(3)$ denotes the rotation group.
@@
In contrast, the composition of Lorentz transformations is more intricate. The key distinction between the Galilean and Lorentz groups lies in their structure: the Lorentz group is simple, meaning it does not naturally decompose into a direct product of rotations and boosts. Unlike in the Galilean case, $O(3)$ is not a normal subgroup of the Lorentz group $O(1,3)$, and the set of Lorentz boosts does not even form a group. The theory of the Galilean transformation composition law is straightforward and well understood. In contrast, the analogous theory of the Lorentz transformation composition law is unsatisfactory.
We conclude that the standard analysis of the relationship between inertial frames overlooks the fundamental difference between Newtonian space and time and the pseudo-Euclidean geometry of space-time. In classical mechanics, the components of the velocity three-vector transform under a Galilean transformation in the same way as the components of the position three-vector $\vec{x}$. This is because, according to classical kinematics, all inertial frames share the same ordinary three-dimensional space. In contrast, in special relativity, different inertial frames have distinct three-dimensional spaces within the framework of space-time. Therefore, unexpected results may arise in special relativity when transitioning from the four-vector formalism back to the three-dimensional velocity vector $\vec{v}$, which is expressed in terms of the four-vector components $(x_{0},\vec{x})$ as $\vec{v}=d\vec{x}/dx_{0}$.
Most texts on special relativity present the relativistic velocity addition formula only for parallel velocities. In this simplified case, Einstein’s velocity addition is both commutative and associative, making it a commutative group operation. However, in the general case, Einstein’s velocity addition $\oplus$ does not form a group.
The relativistic reciprocity principle states that if an inertial frame $K^{\prime}$ moves with velocity $\vec{v}$ relative to another inertial frame $K$, then, reciprocally, the velocity of $K$ relative to $K^{\prime}$ is $-\vec{v}$. This principle suggests that the inverse of a velocity $\vec{v}$ is simply its negation: $\vec{v}^{-1}=-\vec{v}$, analogous to the Galilean addition of velocities. For composite velocities, $\vec{u}\oplus\vec{v}$, reciprocity implies $(\vec{u}\oplus\vec{v})^{-1}=-(\vec{u}\oplus\vec{v})=(-\vec{v})\oplus(-\vec{u})=-(\vec{v}\oplus\vec{u})$. However, since Einstein’s velocity addition $\oplus$ is generally non-commutative, this leads to a contradiction, known as the reciprocity paradox. This paradox is closely related to the Mocanu paradox, as discussed by Mocanu [66] and was latter resolved by Ungar [67] who demonstrated the crucial role of the Wigner (Thomas) rotation in the transformation. Furthermore, Ungar [67] also showed that Einstein’s velocity addition is also non-associative meaning that $(\vec{w}\oplus(\vec{v}\oplus\vec{u}))\neq((\vec{w}\oplus\vec{v})\oplus\vec{u})$. Being non-associative, the Einstein’s addition $\oplus$ is not a group operation. Non-associativity is paradoxical: for a system of four bodies the $\oplus$ of three non-parallel velocities gives two distinct velocities between two bodies.
Most textbooks present the Lorentz transformation as being generated by a boost vector $\vec{v}$, interpreted as the unique reciprocal velocity between two inertial frames. However, we emphasize that the reciprocal property $\vec{v}^{-1}=-\vec{v}$ is inconsistent with Einstein’s two postulates of special relativity  [^55] For a more in-depth discussion on the relationship between the reciprocity axiom and Einstein’s two postulates of special relativity, we recommend the chapter ”How Do You Add Relative Velocities?” by Oziewicz in [69].. The Lorentz isometry group $\Lambda$ acts on four-dimensional spacetime, and all subgroups of the Lorentz group operate on four-vectors in spacetime. Consequently, there is no concept of common (instantaneous) spatial frame within the framework of the Lorentz group. The Lorentz transformation $\Lambda$ does not naturally decompose into a unique combination of a rotation and a boost. Any such decomposition would imply the existence of a preferred frame, contradicting the principle of relativity.
In Minkowski space, where spacetime follows pseudo-Euclidean geometry, equivalence of all inertial frames is not possible. This suggests the existence of an explicit, preferred (initial) inertial reference frame within special relativity, challenging the conventional interpretation. The relationship between inertial frames are not determined solely by their relative velocities 565656 In particular, the presence of non-zero relative velocity cannot, by itself, be a cause of time dilation (see Section 12.3 for more detail)..
The inverse of a velocity vector, $\vec{v}^{-1}=-\vec{v}$, is often presented in textbooks as a self-evident fact requiring no further discussion. However, this assumption contradicts a fundamental axiom of relativity theory—namely, the axiom of pseudo-Euclidean spacetime geometry.
@@
For example, in his contemporary textbook, Barut defines Lorentz boost as (see Eq. (1.28) and the accompanying discussion in [6]):
@@
It is evident that all textbooks, including [6], define the Lorentz boost as a 4x4 basis-dependent matrix. Next we shall discuss some physical applications of the transformation $L_{\vec{v}}$.
@@
Consider a coordinate system $K_{1}$ moving with respect to $K_{0}$ with a velocity $\vec{v}_{01}$ and a frame $K_{2}$ moving with respect to $K_{1}$ with a velocity $\vec{v}_{12}$. It is important to note that $\vec{v}_{01}$ represents the velocity of frame $K_{1}$ as measured by an observer at rest in $K_{0}$, while $\vec{v}_{12}$ represents the velocity of $K_{2}$ as measured by an observer in $K_{1}$. The question then arises: what is the Lorentz transformation from $K_{0}$ to $K_{2}$? Let the corresponding Lorentz transformations be denoted by $L_{1}$, $L_{2}$ and $L$, respectively. Then, according to standard textbooks conventions, we have $L=L_{2}L_{1}$.
@@
Now, consider the case, where $\vec{v}_{01}$ and $\vec{v}_{12}$ are perpendicular. The velocity $\vec{v}_{02}$ of frame $K_{2}$ as observed from $K_{0}$ is given by $\vec{v}_{02}=\vec{v}_{01}+\vec{v}_{12}\sqrt{1-v_{01}^{2}/c^{2}}$. According to reciprocity principle, accepted in all textbooks, an observer in $K_{2}$ sees frame $K_{1}$ moving with velocity $-\vec{v}_{12}$, while an observer in $K_{1}$ sees frame $K_{0}$ moving with velocity $-\vec{v}_{01}$. Similarly, applying Einstein’s velocity addition, we find that the velocity $\vec{v}_{20}$ of frame $K_{0}$ as observed from $K_{2}$ is: $\vec{v}_{20}=-\vec{v}_{12}-\vec{v}_{01}\sqrt{1-v_{12}^{2}/c^{2}}$. According to textbooks, $\vec{v}_{20}$ should be reciprocal of $\vec{v}_{02}$, but they point in different directions. The angle between $\vec{v}_{02}$ and $\vec{v}_{20}$ is the Wigner rotation angle. This reciprocity discrepancy leads to the Mocanu contradiction and highlights the non-associativity of Einstein’s velocity addition.
While composite velocities in Galilean relativity fully determine the relationships between inertial frames, their relativistic counterparts do not. In addition to the relativistic velocity composition between inertial frames, one must also account for the associated Wigner (Thomas) rotations.
In his papers and monographs—written at an advanced mathematical level — Ungar attempted to resolve the non-associativity paradox of Einstein’s velocity addition $\oplus$ by refining the composition law (see [68] and references therein). The improved composition law replaces $\vec{u}\oplus\vec{v}$ with $[\vec{u}\oplus\vec{v},\mathrm{W}(\vec{u},\vec{v})]$, where $\mathrm{W}(\vec{u},\vec{v})$ represents the Wigner (Thomas) rotation generated by $\vec{u}$ and $\vec{v}$. Notably, the binary operation $\oplus$ remains the standard Einstein velocity addition.
However, this approach is unsatisfactory even from a mathematical standpoint for a fundamental reason. As demonstrated in Section 16.6, the Wigner rotation is not merely an additive correction but is intrinsically woven into the structure of relativistic velocity addition itself. Consider the asymptotic case of two perpendicular velocities: $\vec{v}_{01}$ approaching the speed of light and $\vec{v}_{12}$ being small. In this limit, $\gamma=(1-v_{01}^{2}/c^{2})^{-1/2}\gg 1\gg v_{12}/c$. The frame $K_{1}$ with the moving frame $K_{2}$ seems to rotate about frame $K_{0}$. We want now to look at the consequence of this Wigner rotation. In the Section 16.6 we found that in Lorentz coordinatization, the velocity $v_{01}$ of frame $K_{2}$ in frame $K_{0}$, after a boost with velocity $\vec{v}_{12}$, undergoes a simple rotation by an angle $v_{12}/(\gamma v_{01})=v_{12}/(\gamma c)$ without any change in its magnitude. In contrast, Ungar’s approach—based on the standard Einstein velocity addition within an improved composition law—suggests that the total speed of $K_{2}$ in the Lorentz frame $K_{0}$ increases from $v_{01}$ to $v_{01}+v_{12}^{2}/(2\gamma^{2}c)$.
In all special relativity textbooks, the principle of reciprocity states that the inverse of velocity addition must be identical to that in Galilean absolute time. Consequently, Einstein’s velocity addition is based on the relation $dt^{\prime}=dt/\gamma$, which implicitly assumes an absolute time convention. However, it does not account for the interplay between transverse positions and time, which is inherent in Lorentz boost matrices. In reality, every Lorentz boost represents a transformation between reference frames within a Lorentzian framework (i.e., under Lorentz coordinatization). Thus, we conclude that the commonly accepted derivation of velocity composition does not follow from the composition of Lorentz boosts, $L=L_{2}L_{1}$.
@@
Under Lorentz coordinatization, moving frames inherently possess different simultaneity relations. As a result, the velocity $\vec{v}$ and its inverse $\vec{v}^{-1}$ are each tangent to a different instantaneous space and these two spaces are not parallel. Consequently, the inverse velocity cannot be truly reciprocal. The reciprocity of frames contradicts relativity of simultaneity. In Enstein’s velocity addition all velocities $\vec{v}_{01}$, $\vec{v}_{12}$, $\vec{v}_{02}$, are implicitly assumed to lie within the same preferred 3-space as seen by a preferred observer. In other words, the reciprocity of relative velocities holds only under an absolute time framework.
The standard textbook presentation of the relativistic transformations implicitly assumes that observers in $K_{0}$, $K_{1}$, and $K_{2}$ inertial frames share a common 3-space. It is often assumed that the acceleration of a rigid reference frame grid has a direct physical meaning. However, this is a misconception. Different inertial frames have distinct ordinary spaces and implicit assumption of velocity reciprocity contradicts Einstein’s two postulates.
The concepts of observer independence and Lorentz covariance are fundamentally distinct. We conclude that there is a striking conflict between the Principle of Relativity (i.e., the equivalence of all inertial frames) and the axiom of pseudo-Euclidean spacetime geometry. In other words, only one of these principles can hold. Based on experimental evidence, we adopt the Lorentz covariance principle  [^57] However, an alternative perspective exists in the literature. Oziewicz, for instance, favors the Principle of Relativity [69]..
This is a good point to make some general remarks about Einstein’s velocity addition. It is important to recognize that it does not directly follow from the Lorentz transformations. While Lorentz boosts intertwine space and time coordinates, Einstein’s velocity addition relies on the assumption of a shared three-dimensional space across all inertial frames.
Despite this, Einstein’s velocity addition is widely used in relativistic dynamics and electrodynamics, raising a reasonable question: How can this be? After all, we know that this form of velocity addition does not constitute a mathematical group—it is non-associative, and without associativity, the development of a consistent and complete theory would seem impossible.
Earlier, we demonstrated that the main issue lies in the definition of relative velocity. If we postulate, as Einstein did, the principle of relativistic reciprocity, we inevitably encounter the non-associativity of velocity addition. Nevertheless, one should recognize the existence of a preferred (initial) inertial frame without an accelerational history relative to the fixed stars. Based on experimental evidence, we adopt the Sun-centered frame as the preferred inertial frame (see Chapter 15 for further discussion). Once a preferred frame is established, there is one—and only one—well-defined relative velocity between any pair of inertial frames.
Let us apply this approach to an example and observe how it performs in practice. Suppose an observer in the initial inertial frame $K_{0}$ establishes a coordinate system using Einstein’s synchronization procedure, based on light signals emitted by a stationary dipole source, under the assumption that light propagates isotropically at speed $c$. Now consider two additional moving inertial frames, $K_{1}$ and $K_{2}$, whose motions are analyzed by the observer in $K_{0}$. The simplest synchronization method involves maintaining the same set of synchronized clocks, thus preserving simultaneity and adhering to an absolute-time coordinatization in $K_{0}$.
We assume that system $K_{1}$ moves with velocity $v_{x}$ along the $x$-axis relative to $K_{0}$, and $K_{2}$ moves with velocity $v_{y}$ along the $y$-axis relative to $K_{1}$. It is important to note that $v_{x}$ and $v_{y}$ represent the velocities as measured by the observer in $K_{0}$.
Applying Einstein’s velocity addition, the observer in $K_{0}$ concludes that an observer in $K_{1}$ sees frame $K_{2}$ moving with a velocity $v_{y}/\sqrt{1-v_{x}^{2}/c^{2}}$. According to Einstein, the velocity of frame $K_{2}$ as observed directly from $K_{0}$ is given by $\vec{v}_{02}=v_{x}\vec{e}_{x}+v_{y}\vec{e}_{y}$ consistent with Galilean velocity addition within the initial inertial frame.
Now, consider a different perspective. The observer in $K_{0}$ sees the relative velocity of $K_{2}$ with respect to $K_{1}$ as $-v_{y}$. Similarly, an observer in $K_{2}$ sees $K_{1}$ moving with velocity $-v_{y}/\sqrt{1-v_{02}^{2}/c^{2}}$. According to Einstein’s velocity addition, the velocity of $K_{1}$ as observed from $K_{0}$ is $\vec{v}_{01}=v_{x}\vec{e}_{x}$ as expected.
We thus conclude that Einstein’s velocity addition is consistent with Galilean velocity addition when viewed from the initial inertial frame. In this scheme, reciprocity is absent, and all paradoxes are eliminated. Indeed, there is no reciprocity in the relative velocities between $K_{1}$ and $K_{2}$, because time dilation is absolute, not relative, in the presence of a preferred inertial frame.
Now, let us consider the Earth-based frame. For an observer on Earth, the velocity relative to the initial inertial frame is approximately $v_{orb}\simeq 30~{}\mathrm{km/s}$, allowing us to neglect terms of order $v_{orb}^{2}/c^{2}$. As previously demonstrated, current experimental techniques usually lack the sensitivity required to detect the influence of orbital velocity on the relativistic particle dynamics. Consequently, in accelerator physics, the Earth-based frame can be treated with good accuracy as the initial inertial frame.
@@
It is important to emphasize that Einstein’s velocity addition is fully compatible with conventional particle tracking within a single inertial frame, which itself relies on the absolute time convention. Crucially, Einstein’s velocity addition inherently incorporates relativistic time dilation. Similarly, conventional particle tracking also accounts for relativistic time dilation, albeit implicitly, through the use of relativistic mass for moving particles.

### v12

@@ 17.1 Parametrization of Lorentz Transformations
Consider a four-vector, $x=(x_{0},\vec{x})$, with squared-length $x^{2}=\eta_{ij}x^{i}x^{j}=(x_{0})^{2}-|\vec{x}|^{2}$ where $\eta_{ij}=\mathrm{diag}[1,-1,-1,-1]$ is the Minkowski space-time metric. Since Lorentz transformations preserve the space-time interval, they satisfy $x^{\prime}=\Lambda x$ such that $\eta_{ij}=\Lambda^{m}_{i}\Lambda^{n}_{j}\eta_{mn}$. This condition defines the most general $4\times 4$ Lorentz transformation matrix $\Lambda$. The set of all such matrices forms the symmetry group of the Minkowski metric under matrix multiplication, denoted by $O(1,3)$. In particular, $O(1,3)$ is a Lie group known as the Lorentz group.
@@
It is widely accepted that the relationship between inertial frames is fully determined by their relative velocities and orientations. Consequently, a general Lorentz transformation, denoted $\Lambda(\vec{v},\vec{\theta})$, is parameterized by a velocity vector $\vec{v}$ and a rotation vector $\vec{\theta}$, representing a combined boost and spatial rotation. Here, $\vec{v}$ is a vector in three-dimensional Euclidean space, while $\vec{\theta}$ belongs to the Lie algebra of the rotation group $O(3)$, which consists of all possible spatial rotations in three dimensions. This parameterization of the Lorentz group is standard in the literature.
Now, we turn to a significant conceptual and structural challenge associated with this parametrization of the Lorentz group. A comparison with the Galilean group provides valuable insight.
@@
It is well known that the Galilean transformation $G(\vec{v},\vec{\theta})$, which satisfies the group composition law, is isomorphic to the semi-direct product group $\mathbb{R}^{3}\rtimes O(3)$, where $\mathbb{R}^{3}$ represents the group of Galilean boosts and $O(3)$ denotes the rotation group. In this structure, $O(3)$ acts as a normal subgroup, and the composition of transformations remains straightforward.
@@
In contrast, the composition of Lorentz transformations is significantly more intricate. The essential distinction between the Galilean and Lorentz groups lies in their group-theoretic structure: the Lorentz group is a **simple** Lie group, meaning it does not admit a nontrivial decomposition into a direct or semi-direct product of boosts and rotations. In particular, $O(3)$ is not a normal subgroup of the Lorentz group $O(1,3)$. As a result, the combined action of Lorentz boosts and spatial rotations does not obey the same algebraic structure as in the Galilean case.
This leads to an important conclusion: standard textbook analyses of inertial frames often overlook the fundamental geometric difference between Newtonian space and time, and the pseudo-Euclidean geometry of Minkowski spacetime.
In classical mechanics, the components of the velocity three-vector transform under Galilean transformations in the same way as the components of the position vector $\vec{x}$. This correspondence arises because, in Newtonian physics, all inertial observers share a common, absolute three-dimensional space. Time is also absolute, so the spatial position and velocity vectors can be consistently defined across all inertial frames.
However, in special relativity, this intuition fails. When transitioning from the four-dimensional spacetime formalism back to the three-dimensional velocity vector $\vec{v}$, defined as $\vec{v}=d\vec{x}/dx^{0}$, subtle but profound issues emerge. The Lorentz group $O(1,3)$ acts on four-dimensional spacetime, and all of its subgroups also act on four-vectors, not directly on spatial vectors. Therefore, there is no concept of a universal or common instantaneous three-dimensional spatial frame among all inertial observers within the framework of the Lorentz group.
This absence of a shared simultaneity structure underlies many of the non-intuitive features of relativistic kinematics, including the non-associativity and non-commutativity of velocity addition.
Every Lorentz transformation that relates inertial observers corresponds to an isometry of the Minkowski metric. However, when two reference frames are related by a general Lorentz transformation—i.e., an element of the Lorentz group O(1,3)O(1,3)—the concept of relative velocity between them is not immediately well-defined.
To extract the relative velocity, one typically attempts to factor the Lorentz transformation into a pure boost (parameterized by a velocity vector) and a spatial rotation. However, such a factorization is not unique: there are many ways to choose the rotation subgroup $O(3)$ within $O(1,3)$, and different choices can lead to different definitions of the relative velocity $\vec{v}$.
Mathematically, the space of all boosts (i.e., all Lorentz transformations modulo spatial rotations) is the coset space $O(1,3)/O(3)$. Since $O(3)$ is not a normal subgroup of $O(1,3)$, different embeddings of $O(3)$ yield different decompositions and thus different velocities. The commonly used decomposition, often referred to in the literature as the ”standard choice,” implicitly assumes a privileged reference frame—a specific frame in which the rotation subgroup $O(3)$ is defined.
In the absence of such a preferred frame, the Lorentz transformation between two systems remains well-defined as an isometry of spacetime, but the notion of a unique relative velocity becomes ambiguous. To define a single, well-determined velocity, one must fix a specific reference frame as distinguished. Physically, this can be modeled by introducing a preferred timelike vector field on spacetime, which selects a particular congruence of worldlines—effectively, a material or ”ether-like” reference frame.
However, this dependence on a distinguished timelike direction violates the relativity principle, which asserts the equivalence of all inertial frames. Thus, within the geometry of pseudo-Euclidean spacetime, the equivalence of all reference systems cannot be maintained.
17.2 Observer-Independence and Lorentz Covariance
A standard approach to representing Lorentz transformations is via transformation matrices acting on spacetime four-vectors. Any Lorentz transformation can, in principle, be decomposed into a product of a rotation and a boost. However, the explicit form of this decomposition depends on the choice of reference frame, which in the matrix formalism is determined by the selection of a distinguished timelike vector, typically denoted by $e_{0}$, in a chosen basis.
@@ 17.2 Observer-Independence and Lorentz Covariance
For example, in his widely used textbook, Barut defines a Lorentz boost (see Eq. (1.28) and the accompanying discussion in [66]) as:
@@
where $\beta=v/c$ and $\gamma=1/\sqrt{1-\beta^{2}}$. This boost corresponds to motion along the $x$-axis and is defined relative to a specific basis.
However, the Lorentz boost is not fully determined by the velocity parameter $\vec{v}$ alone. The very notion of velocity presupposes a distinction between space and time—i.e., a $1+3$ decomposition of spacetime—which requires the selection of a specific timelike direction. This decomposition identifies the observer through the choice of a basis vector $e_{0}$, which represents the time direction in that observer’s rest frame.
As made explicit in Barut and other standard references [66], the Lorentz boost is expressed as a $4\times 4$ matrix relative to a given orthonormal basis. In the “standard” formulation, the preferred inertial frame is associated with the basis vector $e_{0}=(1,0,0,0)$. Rotations then occur within the three-dimensional spatial subspace orthogonal to $e_{0}$, and any boost involving a velocity vector $\vec{v}$ orthogonal to $e_{0}$ implicitly defines the boost with respect to this preferred frame. Thus, even though Lorentz transformations are covariant under changes of inertial observers, their explicit matrix representations—and the definition of velocity itself—rely on a choice of observer.
Decompositions of spacetime based on different choices of the timelike basis vector $e_{0}$ are mutually incomparable, as they correspond to distinct three-dimensional spatial slices for which no canonical identification exists.
Successive Lorentz boosts are composed through matrix multiplication, assuming both boosts are expressed in the same basis. Specifically, if $L_{\vec{v}_{1}}$ is the boost from frame $K_{0}$ to frame $K_{1}$, and $L_{\vec{v}_{2}}$ is the boost from $K_{1}$ to $K_{2}$, then their composition is given by: $L_{\vec{v}_{2}}L_{\vec{v}_{1}}$. Here, $\vec{v}_{1}$ is the velocity of frame $K_{1}$ relative to the preferred frame $K_{0}$, and $\vec{v}_{2}$ is the velocity of frame $K_{2}$ relative to $K_{1}$, as observed from $K_{0}$.
It is crucial to distinguish between observer-independence and Lorentz covariance. Although Lorentz covariance refers to the invariance of physical laws under Lorentz transformations, it does not imply that all inertial frames are physically equivalent in a pseudo-Euclidean spacetime. The geometry of Minkowski space allows for Lorentz covariance without guaranteeing the equivalence of all inertial observers. This subtle distinction opens the door to the possibility of a physically preferred inertial frame, thereby challenging the traditional interpretation of the principle of relativity.
In this context, relative velocity alone does not fully characterize the relationship between inertial frames.  [^55] In particular, a non-zero relative velocity does not, by itself, cause time dilation. See Section 12.3 for further discussion. A deeper tension emerges between the principle of relativity—which asserts the equivalence of all inertial frames—and the geometric structure of Minkowski spacetime. In fact, only one of these principles can hold unconditionally.  [^56] For a more detailed analysis, see “How Do You Add Relative Velocities?” by Oziewicz in [67]. Oziewicz argues that within pseudo-Euclidean spacetime, the equivalence of all reference frames is not possible.
Given the strength of experimental evidence supporting Lorentz invariance, we adopt Lorentz covariance as the foundational principle of the theory.
As discussed in earlier chapters, further mathematical considerations support the view that pseudo-Euclidean geometry does not entail the complete equivalence of all inertial frames. One key insight comes from analyzing the continuity of the metric tensor under coordinate transformations. The Langevin metric, for instance, arises from a continuous deformation of the Minkowski metric when transitioning between an inertial frame and an accelerated frame. This smooth transition can be modeled via a Galilean boost relating the coordinate systems of the inertial and accelerated observers.
However, restoring the diagonal Minkowski form in the accelerated frame often requires an abrupt coordinate transformation, resulting in a discontinuity in the metric tensor. This discontinuity signals a breakdown in the symmetry between inertial frames: if a Lorentz transformation fails to preserve the smoothness of the metric, then the equivalence of the frames involved is compromised.
This raises a natural and important question: How can observers identify a preferred inertial frame?
Within the structure of the Lorentz group, relative velocity between frames enters exclusively through the boost transformations. However, a Lorentz boost is always defined relative to a specific reference frame—typically one associated with a distinguished timelike vector. Special relativity, in its standard formulation, does not specify such a frame; yet physical observations suggest the existence of a unique inertial frame—one not subjected to prior accelerations, such as relative to the fixed stars.
There is, however, no intrinsic contradiction between the mathematical framework of special relativity and the physical asymmetry introduced by selecting a preferred frame. The theory is built on the requirement that physical laws remain invariant under Lorentz transformations. But special relativity is not a self-contained theory in the sense of fully determining physical phenomena from dynamical laws alone—initial conditions must also be specified. The principle of relativity asserts that the form of physical laws is independent of velocity, but not necessarily that initial conditions must be velocity-independent.
While Lorentz coordinatization appears to establish symmetry among inertial frames, the asymmetry emerges from the history of acceleration. Once a frame has undergone acceleration, information about that acceleration is not encoded in the transformation laws themselves but in the initial conditions from which they operate.
Consider, for instance, the relativistic aberration of particles emitted from a source at rest in an accelerated frame. As discussed in Chapter 9, special relativity predicts that—alongside a static electric field—a magnetic field also arises in the accelerated frame, leading to deviations in the observed trajectory of particles. This deviation is a measurable consequence of the frame’s prior acceleration.
The mathematical framework of special relativity remains entirely self-consistent when formulated relative to an initial inertial frame—one that has not undergone prior acceleration with respect to the distant stars. Such a frame functions as a *privileged reference system*. As demonstrated in Chapter 7, when the laws of physics are formulated relative to this frame, they accurately account for all observed phenomena, including those reported by accelerated observers.
In particular, phenomena like the aberration of light in an accelerated frame can be fully derived using standard tools of special relativity—specifically, Lorentz transformations and Wigner rotations—when analyzed from the standpoint of the initial inertial frame.
17.3 Einstein Velocity Addition and Conventional Particle Tracking
Composition of relativistic velocities, also named as ”Einstein’s addition of velocities” is somewhat confusing due to fact that this addition is nonassociative. This is contrast with associativity of the Lorentz group. Most texts on special relativity present the relativistic velocity addition formula only for parallel velocities. In this simplified case, Einstein’s velocity addition $(\oplus)$ is both commutative and associative, making it a commutative group operation analogous to the Galilean velocities addition $(+)$. However, in the general case, Einstein’s velocity addition $\oplus$ does not form a group.
The relativistic reciprocity principle states that if an inertial frame $K^{\prime}$ moves with velocity $\vec{v}$ relative to another inertial frame $K$, then, reciprocally, the velocity of $K$ relative to $K^{\prime}$ is $\vec{v}$. This principle suggests that the inverse of a velocity $\vec{v}$ is simply its negation: $\vec{v}^{-1}=-\vec{v}$, analogous to the Galilean addition of velocities. This is often presented in textbooks as a self-evident fact requiring no further discussion.
For composite velocities, $\vec{u}\oplus\vec{v}$, reciprocity implies $(\vec{u}\oplus\vec{v})^{-1}=-(\vec{u}\oplus\vec{v})=(-\vec{v})\oplus(-\vec{u})=-(\vec{v}\oplus\vec{u})$. However, since Einstein’s velocity addition $\oplus$ is generally non-commutative, this leads to a contradiction, known as the reciprocity paradox. This paradox is closely related to the Mocanu paradox, as discussed by Mocanu [68] and was latter resolved by Ungar [69] who demonstrated the crucial role of the Wigner rotation in the transformation.  [^57] It is sometimes referred to as the Thomas rotation Furthermore, Ungar [69] also discovered that Einstein’s velocity addition is non-associative meaning that $(\vec{w}\oplus(\vec{v}\oplus\vec{u}))\neq((\vec{w}\oplus\vec{v})\oplus\vec{u})$. Being non-associative, the Einstein’s addition $\oplus$ is not a group operation. Non-associativity is paradoxical: for a system of four bodies the $\oplus$ of three non-parallel velocities gives two distinct velocities between two bodies.
@@ 17.3 Einstein Velocity Addition and Conventional Particle Tracking
Let us illustrate non-commutativity of Einstein’s velocity addition using only simple case of relativistically combining perpendicular velocities. Consider a coordinate system $K_{1}$ moving with respect to $K_{0}$ with a velocity $\vec{v}_{01}$ and a frame $K_{2}$ moving with respect to $K_{1}$ with a velocity $\vec{v}_{12}$. It is important to note that $\vec{v}_{01}$ represents the velocity of frame $K_{1}$ as measured by an observer at rest in $K_{0}$, while $\vec{v}_{12}$ represents the velocity of $K_{2}$ as measured by an observer in $K_{1}$. The question then arises: what is the velocity transformation from $K_{0}$ to $K_{2}$?
@@
Now, consider the case, where $\vec{v}_{01}$ and $\vec{v}_{12}$ are perpendicular. The ”perpendicular” means that velocity $\vec{v}_{12}$ is perpendicular to line motion of frame $K_{0}$ in $K_{1}$. The velocity $\vec{v}_{02}$ of frame $K_{2}$ as observed from $K_{0}$ is given by $\vec{v}_{02}=\vec{v}_{01}+\vec{v}_{12}\sqrt{1-v_{01}^{2}/c^{2}}$. According to reciprocity principle, accepted in all textbooks, an observer in $K_{2}$ sees frame $K_{1}$ moving with velocity $\vec{v}_{12}$, while an observer in $K_{1}$ sees frame $K_{0}$ moving with velocity $\vec{v}_{01}$. Similarly, applying Einstein’s velocity addition, we find that the velocity $\vec{v}_{20}$ of frame $K_{0}$ as observed from $K_{2}$ is: $\vec{v}_{20}=-\vec{v}_{12}-\vec{v}_{01}\sqrt{1-v_{12}^{2}/c^{2}}$. According to textbooks, $\vec{v}_{20}$ should be reciprocal of $\vec{v}_{02}$, but they point in different directions. This reciprocity discrepancy leads to the Mocanu contradiction and highlights the non-associativity of Einstein’s velocity addition.
@@
Under Lorentz coordinatization, moving frames inherently possess different simultaneity relations. As a result, the velocity $\vec{v}$ and its inverse $\vec{v}^{-1}$ are each tangent to a different instantaneous space and these two spaces are not parallel. Consequently, the inverse velocity cannot be truly reciprocal. The reciprocity of frames contradicts relativity of simultaneity. The commonly accepted derivation of velocity composition does not follow from the composition of Lorentz boosts and based on absolute time convention. In Enstein’s velocity addition all velocities $\vec{v}_{01}$, $\vec{v}_{12}$, $\vec{v}_{02}$, are implicitly assumed to lie within the same preferred 3-space as seen by a preferred observer. It is often assumed that the acceleration of a rigid reference frame grid has a direct physical meaning. However, this is a misconception.
@@
Despite this, Einstein’s velocity addition is widely used in relativistic dynamics raising a reasonable question: How can this be? It is important to emphasize that Einstein’s velocity addition is fully compatible with conventional particle tracking within the initial inertial frame, which itself relies on the absolute time convention. Crucially, Einstein’s velocity addition inherently incorporates relativistic time dilation. Similarly, conventional particle tracking also accounts for relativistic time dilation, albeit implicitly, through the use of relativistic mass for moving particles.